OpenStudy (gabylovesyou):

Let f(x) = x2 - 8x + 5. Find f(-1). = 14 right ? @blurbendy

OpenStudy (anonymous):

do i need to answer this or no?

OpenStudy (blurbendy):

yes, correct

OpenStudy (hari5719):

it is right

OpenStudy (anonymous):

This is how > :basically just insert -1 wherever there is an x. (I'm assuming here that x2 is x squared) f(-1) = (-1)^2- 8(-1)+5 = 1+8+5 = 14 Any negative multiplied by a negative will give you a positive. So... -1(-1)= 1 and -8(-1)=8

OpenStudy (anonymous):

Well NEVERMIND

OpenStudy (anonymous):

-1^2 - 8(-1) + 5 1 - 8 + 5 = 12 Answer is 12

OpenStudy (hari5719):

GUYS VIEW THIS FOR THE ANSWER http://www.wolframalpha.com/input/?i=f%28x%29+%3D+x2+-+8x+%2B+5.+Find+f%28-1%29.

OpenStudy (anonymous):

\( f(-1) = -1^2 - 8(-1) + 5 \) \( f(-1) = 1 - 8 + 5 = 12\) \( f(-1) = 12\) Answer is 12

OpenStudy (anonymous):

It is 14 only if that is not a x^2 but x2, which I am thinking it is meant to be a quadratic.

OpenStudy (hari5719):

@Gabylovesyou THe answer is 14 u are right proof http://www.wolframalpha.com/input/?i=f%28x%29+%3D+x2+-+8x+%2B+5.+Find+f%28-1%29.